How many moles of carbon atoms are needed to combine with 4.87 mol Cl to form the substance C2Cl6?
1
Expert's answer
2020-06-17T05:56:06-0400

Solution.

x 4,87

2C + 3Cl2 = C2Cl6

2 3

Let n(C) = x

x = 4.87*2/3 = 3,247 mol.

n(C) = 3,247 mol.

Answer: n(C) = 3,247 mol.